The kiosk systems have all been replaced by new Dell Optiplex 7010 systems. Please notify the lab staff if you discover
any problems with these systems.
Matlab has been updated to 2013b. For more information on the new release see
the Mathworks web site Matlab 2013b
A replacement color MFP has been installed in D108. More information on the printer and its features
can be found on the CSL printers page.
A new Dell Optiplex 7010 has been install in D108 for use with the new photo scanner. This system has photo and video editing software installed.
An Epson Perfection V600 Photo Scanner has been installed in D108. Please see the information page for details
and operation instructions.
The computer in classrom D106 has been replaced with a Optiplex 7010.
Please contact problem if you encounter any issues.
The three classrooms in the LSRC D106, D243, D344 have received updated projectors.
The new projectors are NEC NP-PA550W-13ZL.
Please let us know if you have any questions about them.
Webmail has been updated to .0.8.1. For more information on the new release
see the CS Webmail website
Mathematica has been updated to 8.0.4. For more information on the new release see
the Mathematica web site Mathematica 8.0.4
Matlab has been updated to 2012b. For more information on the new release
see the Mathworks web site Matlab 2012b
Maple has been updated to version 16. For more information on the new release see
the Maple web site Maple 16
We've upgraded and migrated our CUPS
(Unix printing) server from version 1.1.23 running on Solaris to 1.5.2 on Linux.
Print jobs are now dispatched as PDF natively to printers rather than
The matlab license has been updated.
The linux cluster has been updated to version 12.04 of the Ubuntu operating system.
For more information take the Ubuntu tour
or see the CS cluster page.
Work continues replacing the older solaris/sparc equipment with linux servers.
We have replaced many of the core services already including network monitoring, NIS, and DHCP. 7/07/12
Our central file servers have been upgraded to a clustered file server.
The two FAS3240 servers will serve most of the department's data.
As a result of the expansion we will be raising quotas in the coming weeks. 6/30/12
Matlab has been updated to 2012a. For more information on the new release see the
Mathworks web site Matlab 2012a
Ten new linux cluster machines have been added.
The machines all have 64GB of RAM, and dual,quad-core processors.
Matlab has been updated to 2011b. For more information on the new release see the Mathworks
web site Matlab 2011b
The lectopia system has been replaced with panopto. You can access all of the CS
content at compsci.capture.duke.edu.
Some content is restricted to NetID login. If you have not changed your NetID password
within the last year, you will need to do so in order for the new system to authenticate you.
More information about panopto can be found at
OIT's panopto site.
Instructions for accessing the CS file systems from a Mac have been updated.
You can view the updated instructions here, MacOS X Filesharing.
Our annual machine upgrade will begin shortly. We will be upgrading
any CS supported desktop to the latest release of Ubuntu. Also, we will be
replacing many Optiplex 755 systems with new Optiplex 980 systems. For details
on these new computers, please see the
desktop computer list.
We are replacing the remaining HP 4350 network printers with
HP P4015dn printers. Please see the
for updates. If you experience any issues, please report them
to the lab staff; 8/05/10
Alert: There was a problem with a critical
service on the mail server today. This problem has been corrected and all
services should be functioning properly. Please report any additional
issues you experience to the lab staff; 6/29/10
The machine that houses webmail experienced a power supply
failure yesterday. It has been restored to service. Please contact
the lab staff if you experience any issues; 6/29/10
Note: We will be patching the main web
server on Friday th 11th of June. The downtime may last as long as
30 minutes while the patches are applied. The server should be back
in service by 5am. Please let us know if you have any concerns or
Note: There will be a test of the
fire alarm system
in the building today. This is only a test and you do not need
to leave the building; 6/03/10
Note: Department members should be aware
of possible issues resulting
from the recent OIT migration to Microsoft Exchange; 5/27/10
Note: The main file server
upgrade was successfully completed this morning. Please
contact the lab staff if you experience any
Alert: On Thursday, May 20, the lab
staff will update the main file server for the department,
in the morning between 4:45 and 5:15am. Users may experience some
disruption in accessing their files at this time.; 5/18/10
As a reminder to faculty, due to recent changes in AT&T's usage
policy, we have canceled our AT&T wireless accounts. Please see the
mobile FAQ page for more information;
Alert: Department members sitting in
D-wing perimeter offices should be on the lookout for water leaks
from the radiators and contact
the facilities people if detected; 5/11/10
Notice: OIT will be rebooting our
switches on May 5th between 5-6am. The actual downtime should be less
than 10 minutes, but they have allotted 1 hour. Please let us know
if you experience issues after the update; 5/3/10
Alert: It appears there is a problem
with DNS for external sites this morning. Most popular sites are
working since those addresses are cached in our servers, but sites
accessed less frequently are returning errors. We do not have any
more information on the problem yet; 5/3/10
Matlab has been updated to the latest release (2010a). Release
2010a marks the first release without support for the SPARC
platform. Accordingly, the login machines will not receive the
update. The 2009b release should continue to work for sometime
though. Please report any problems you experience to the lab
Notice:The lab staff will upgrade the
SAMBA server (share) tomorrow at 5am. This should cause no issues, but
please notify us if you experience problems.;
Alert: the issue with the
server housing both share
and the /usr/xtmp/ filesystem has reoccurred. We are
continuing to monitor the situation.; 2/07/10
Alert: There will be a brief downtime
for share tomorrow at 4:45am. It should last no more
than 5 minutes. This is part of an ongoing effort to track down
a software issue causing
intermittent outages.; 2/04/10
All of the machines in the linux cluster
have been updated to use the latest revision of Ubuntu 9.10 (Karmic Koala);
bug in the department's spam-processing
software may have caused legitimate email to be marked as spam; please check
you spam folders; the bug has now been corrected; 1/02/10;
As of today, the department has enabled
Spam filtering by default
on all new user accounts. This is an opt-out system; that is, you may
opt in or out or make adjustments however you like, but initially an
account will be opted in; 8/11/09
Users who forward e-mail to gmail (or other) accounts are advised to
filter spam before forwarding;
failure to do so causes the department server to be tagged in anti-spam
databases and causes forwarded messages to be delayed or rejected; 7/14/09
Notice: The faulty network switch was
successfully replaced last night; 7/02/09
Notice: Last night the lab experienced the
failure of a redundant module in the network switches, starting at about 7pm.
While everything generally appears to have continued working, there might have
been some issues related to the failure. We have disabled the redundant
module and a replacement has been ordered. We will follow up later if the
replacement involves any downtime. Please let us know if you continue to
experience issues.; 7/01/09
Department desktop OS changing to
Ubuntu Linux distribution; 6/09/09
Department-maintained fileserver space set aside for
Time Machine users; 5/26/09
Notice: This morning, the department
web server experienced a problem that caused intermittant
or interupted service to the web site, and to various other
department services. The issue has been resolved and we are
keeping an eye on the web server to make sure there are no
further issues; 5/15/09
Alert: A power
has affected many core OIT services; 5/06/09
Changes are being made to the departmental
accounts maintained for faculty use during travel; 5/06/09
Users should be aware of a problem with the driver for the new Sharp
copiers, where they assume an incorrect paper size. Please follow
as a temporary measure; 4/02/09
The department network equipment will be upgraded on Jan 6, 2009,
please see the announcement
for more details; 12/10/08
The department's USENET server will be
(pending check of actual usage of the service, on October 1, 2008; 9/24/08
The /usr/project/xtmp/ disk storage area has been
upgraded to four terabytes on a RAID-5 w/ spare, and retention time
has been increased to 180 days. For additional information, please see our
Disk Usage Policy page;
A new multifunction color copier was added to room D108. In addition to
printing, you should now be able to scan documents and send them via email
to anyone! But you will need to update your drivers. For information and drivers
see the CS Printers list (See printer pr108c) 9/14/06